OpenStudy (anonymous):

what is the integral of ( x^2+2/X^2-1)dx?

OpenStudy (jamesj):

Well, what's the integral of x^2 ? and of 2/x^2 and of -1?

OpenStudy (jamesj):

So what's the integral of x^2 dx ?

OpenStudy (anonymous):

x^3/3

OpenStudy (jamesj):

right. So now integrate the other two terms.

OpenStudy (anonymous):

the 2x?

OpenStudy (anonymous):

integral 2/x^2 dx and integral -1 dx

OpenStudy (anonymous):

*i don't understand. what is the integral of x^2/x^2-1?

OpenStudy (anonymous):

am i going to use u substitution?

OpenStudy (jamesj):

No. What is the integral of x^(-2) dx ?

OpenStudy (anonymous):

uhm. x^(-1)/-1?

OpenStudy (jamesj):

Yes, that's right. Or just -1/x. We are evaluating the integral one term at a time:\[\int\limits ( x^2+2/x^2-1) \ dx = \int\limits x^2 \ dx + \int\limits 2x^{-2} \ dx + \int\limits -1 \ dx \]

OpenStudy (jamesj):

We now have the first two terms. Can you take it from here?

OpenStudy (anonymous):

ah. i see. thanks. kamsahamnida!

OpenStudy (anonymous):

uhm. wait. why did it become integral of 2x^-2?

OpenStudy (jamesj):

because 1/x^2 = x^(-2)

OpenStudy (anonymous):

yes but it also has +1 doesn't it matter? so you mean the all the terms in the denominator goes to the numerator like (x^2-1)^-1?

OpenStudy (jamesj):

oh, your integral is this: \[\int\limits \left( x^2 + \frac{2}{x^2 - 1} \right) \ dx\] ?

OpenStudy (jamesj):

You've got to use brackets otherwise it's not clear.

OpenStudy (anonymous):

uhm no. the one you wrote first was my integral..

OpenStudy (jamesj):

Ok. So in this case you want to write 2/(x^2 - 1) as partial fractions 2/(x^2 - 1) = A/(x-1) + B/(x+1) Now find A & B. And the reason we do that is the terms on the right-hand side are easy to integrate.

OpenStudy (jamesj):

So it is this: \[\int\limits (x^2 + 2/x^2 - 1) \ dx\]

OpenStudy (jamesj):

Yes?

OpenStudy (anonymous):

yes.

OpenStudy (jamesj):

So ... integrate terms by term x^2 2x^(-2) and -1

OpenStudy (anonymous):

why is it 2x^(-2)?

OpenStudy (jamesj):

because\[\frac{2}{x^2} = 2x^{-2}\]

OpenStudy (jamesj):

because 1/a = a^(-1)

OpenStudy (jamesj):

e.g., 10^(-2) = 1/100

OpenStudy (anonymous):

yes i get that but did the x^2 came from the denominator beacuse the x^2 in the denominator still has -1

OpenStudy (jamesj):

\[\frac{2}{x^2} = 2 \frac{1}{x^2} = 2 (x^2)^{-1} = 2 x^{2(-1)} = 2x^{-2}\]

OpenStudy (jamesj):

1/x = x^(-1) 1/x^2 = x^(-2) 1/x^n = x^(-n)

OpenStudy (anonymous):

can we do that directly? didn't the \[x ^{2}\] have -1?

OpenStudy (jamesj):

The point is \[\int\limits \frac{2}{x^2} \ dx = - \frac{2}{x} \ \ (+C)\] however you arrive at that result.

OpenStudy (anonymous):

thanks. uhm last question how do you integrate (x^2/x^2-1)

OpenStudy (jamesj):

\[\frac{x^2}{x^2 - 1} = 1 + \frac{1}{x^2 -1} = 1 + \frac{A}{x-1} + \frac{B}{x+1}\] This is called a partial fraction decomposition. You need now to find the A and B which make this equation correct and then integrate every term on the RHS.

OpenStudy (anonymous):

OMG.! that looks like its difficult.

OpenStudy (anonymous):

wait. my integral was [(x^2+2)/(x^2-1)]dx so its the same with ∫(x2+2/x2−1) dx. the answer is −2x (+C)?

OpenStudy (jamesj):

Then \[\frac{x^2 + 2}{x^2 - 1} = 1 + \frac{3}{x^2 - 1}\] Now use the partial fractions technique above and if you don't know it, go to your text book. It's really not so bad, but it requires familiarity.